# From Patchwork to Network: A Comprehensive Framework for Demand Analysis and Fleet Optimization of Urban Air Mobility

### v1

@@ 1 Introduction
Urban transportation systems around the world face unprecedented strain due to the increase in population density and the expansion of the urban landscape [1]. Recent studies indicate that approximately 56 58% of the world’s population live in urban areas [2, 3], leading to severe traffic congestion, economic losses, and environmental degradation, particularly in megacities such as New York, Los Angeles and Tokyo [4]. Addressing these challenges requires prioritizing strategic planning of transportation networks and requires innovative approaches to urban mobility.
In response, urban air mobility (UAM) has emerged as a transformative paradigm, proposing the use of low-altitude airspace for on-demand passenger transport [5]. The development of electric vertical take-off and landing (eVTOL) aircraft by pioneering companies such as Joby Aviation, EHang, Archer Aviation, and Volocopter is a critical enabler of this vision, promising a safer, quieter, and more sustainable alternative to ground-based travel [6]. However, the immense capital investment required to build a new network from the ground up presents a significant barrier [1, 7]. This challenge highlights the strategic importance of existing aviation infrastructure. For example, the New York and Los Angeles metropolitan areas contain 25 and 24 regional airports, respectively, many with surplus capacity [8, 7]. Therefore, re-purposing existing and underutilized regional airports offers a viable entry point for initiating UAM operations.
However, transitioning UAM from a promising concept to a viable reality presents immense system-level challenges beyond infrastructures. Successful launch and operation of an UAM service requires solving a series of additional problems, including demand modeling and estimation, fleet sizing, network and corridor design, and real-time fleet management under uncertainty[6]. Among these, heterogeneous fleet composition emerges as a critical factor, as it directly influences capital investment and operating efficiency [9, 10]. Optimizing aircraft selection to match spatiotemporal demand variations minimizes seat wastage while ensuring service coverage, thereby crucially influencing system viability.
Furthermore, conventional transportation analysis tools are often inadequate for this task [11]. They struggle with the scale of regional analysis and often lack the ability to model the dynamic feedback loop between emerging air traffic and existing ground transportation systems. The significant computational time required by conventional simulators creates a performance bottleneck, hindering the large-scale, iterative analysis needed for robust system design and optimization [12, 13].
@@ 2.1 UAM Demand Calculation
Strategic network planning is the cornerstone of successful implementation of UAM, and it fundamentally relies on accurate travel demand forecasting [5, 6, 14]. This section reviews existing literature on UAM demand estimation to highlight a critical methodological limitation: the predominant use of static models that neglect the dynamic feedback between air and ground transportation systems.
Current research primarily estimates UAM demand by modeling mode choice based on the attributes of trips across different transportation options. In this framework, a trip qualifies for UAM if it provides substantial savings in travel time or cost relative to ground transportation [15, 1, 16] Some studies also suggest that airport access and egress times influence demand [17, 18, 14]. Researchers mainly draw on static data to perform demand analysis, including current airport location [14, 19], Uber trip data [19], census-level socioeconomic data [20, 21], cell phone data[20], and ground-level transportation data [22, 15]. Different application of UAM also impact the forecasting of demand, a study in Milan found UAM airport shuttles were more attractive than on-demand air taxis, capturing a potential modal share of 2-5% for airport trips, compared to 1-3% for air taxis [17].
The critical limitation of this static approach is its inability to account for the dynamic feedback loop between UAM and ground transportation systems [23]. As UAM adoption increases, it diverts traffic from terrestrial routes, which can partially alleviate road congestion. By analyzing demand based on a fixed, pre-existing state of congestion, static models overlook this shifting equilibrium. As a result, they are likely to produce an overestimation of the long-term sustainable demand for UAM services. Future research must therefore evolve towards dynamic models that can capture the reciprocal effects between air and ground mobility to generate more realistic and reliable demand forecasts [9].
@@ 2.2 UAM Operation Simulator
The operational management of UAM is exceptionally complex, involving dynamic passenger demand, intricate airspace interactions, and energy-constrained vehicle fleets [24, 25, 26]. Due to these interconnected complexities, analytical solutions are often intractable, making high-fidelity simulation platforms indispensable as virtual testbeds for developing and validating operational algorithms [27, 28].
@@
These simulators provide a dynamic environment to assess the performance and robustness of various management strategies before real-world deployment. For instance, in the domain of on-demand fleet scheduling, simulators are used to evaluate the effectiveness of optimization methods like Mixed-Integer Linear Programming (MILP) and imitation learning in minimizing passenger wait times [29, 30, 31]. Simulators also provide the necessary environment to assess the practical viability and effectiveness of heuristic approaches, such as Genetic Algorithms (GA) and Particle Swarm Optimization (PSO), for managing large-scale, heterogeneous fleets [9].
@@
Furthermore, simulators are critical for validating advanced airspace management strategies. To ensure safety and efficiency in high-density airspace, Demand-Capacity Balancing (DCB) has emerged as another focal point, with solutions relying on advanced AI techniques like Multi-Agent Reinforcement Learning (MARL) for strategic traffic deconfliction [32, 33, 34] and AI-driven urban airspace monitoring method [35]. At the level of real-time dispatching, VertiSync, a centralized, conflict-free takeoff scheduling policy that explicitly integrates both trip request servicing and critical vehicle rebalancing within UAM networks [36]. Building on this, our previous research introduced a Multi-modal Regional-scale Traffic Simulation (MRTS), which provides an integrated environment to model the coupled dynamics of passengers, aircraft, and ground traffic, enabling more holistic algorithm validation [23, 10].
@@
While these operational algorithms are powerful in theory, their practical development and large-scale validation are constrained by the performance of the simulation environments [11, 37, 13]. For any operational strategy, a simulator must validate thousands of diverse, large-scale vehicle scenarios. As a result, the simulator’s speed directly determines the comprehensiveness and efficiency of this validation process [38, 39]. Further more, this constraint is even more acute for AI-driven methods like reinforcement learning, where the simulator is not just a validation tool but an integral part of the training loop [33].
@@
Presently, the vast majority of UAM simulation platforms are built on CPU-based architectures. This creates a computational bottleneck, rendering large-scale, real-time simulations computationally prohibitive and thus severely limiting the scope and speed of algorithmic validation [40]. Therefore, a critical research gap exists between the needs of advanced algorithms and the capabilities of current simulators. Our previous study has developed LPSim, a Large-Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework, which can help boost the computation process [12].
@@ 3.1 UAM and Ground Trip Simulation
Our simulation’s primary objective is to seamlessly integrate UAM into existing transportation systems as a multimodal mode. This integration encompasses the entire journey, from first-mile to last-mile ground segments to the air trips. Consequently, our framework explicitly captures the intricate interplay between UAM operations and traditional ground transportation [12].
@@
5:   Simulate car following, lane changing, and gap acceptance as detailed in [12]
@@ 4.1 Airports and Ground Transportation Data
Urban travel demand data was obtained from the San Francisco Chained Activity Modeling Process-6 (SF-CHAMP) framework, an advanced model developed by the San Francisco County Transportation Authority (SFCTA) that predicts transportation patterns based on infrastructure, travel behaviors, and socio-economic data [41, 42]. The dataset reflects a typical weekday, containing approximately 17.8 million trips over a full day and 6.6 million during the morning period (12:00 AM–12:00 PM), with each trip record detailing attributes such as origin, destination, and transportation mode [43]. Finally, the airport network consists of 21 general aviation airports in the Bay Area, with detailed specifications for each provided in Table 1.
To estimate the potential capacity of each airport in the Bay Area, we evaluated both its ground-side and air-side limits. The maximum ground transportation capacity was calculated using an average vehicle occupancy factor [44]. Concurrently, the maximum air-side capacity was determined by modeling continuous operations at full throughput, assuming a 90-second interval between takeoffs or landings, with the largest suitable aircraft for each facility. The resulting capacity estimates for the airport network are illustrated in Figure 3.
@@ 4.2 Aircraft’s Flight Simulation
For this study, individual flights are modeled using a detailed five-phase profile based on FAA standards [45]. The specific parameters for each phase are defined as follows:
@@ 4.2.3 Travel Time for Aero Trips
To determine the minimum operational interval for each airport runway, we apply the constraints outlined in FAA Order 7110.65 [45]. This interval is a function of two primary components: Runway Occupancy Time (ROT) and the Aircraft Separation Requirement (ASR). For eVTOLs, we assume an ROT of approximately 30 seconds, with a slightly longer ROT for general aviation aircraft like Cessna 208 Caravan. The ASR, which ensures safety between successive aircraft, is set at 60 seconds under Visual Flight Rules (VFR) for small aircraft following another. Therefore, the Minimum Time Interval (T) is the sum of these values, which yields 90 seconds ($30sROT+60sASR$). Consequently, the maximum theoretical capacity for a single runway is 40 operations per hour ($3600s/90s$), where an operation can be either a take-off or landing.

### v2

@@ 1 Introduction
Urban transportation systems around the world face unprecedented strain due to the increase in population density and the expansion of the urban landscape [hae_choi_exploring_2022]. Recent studies indicate that approximately 56 58% of the world’s population live in urban areas [li_rethinking_2024, world_bank_urban_2024], leading to severe traffic congestion, economic losses, and environmental degradation, particularly in megacities such as New York, Los Angeles and Tokyo [qu_demand_2024]. Addressing these challenges requires prioritizing strategic planning of transportation networks and requires innovative approaches to urban mobility.
In response, urban air mobility (UAM) has emerged as a transformative paradigm, proposing the use of low-altitude airspace for on-demand passenger transport [cohen_urban_2021]. The development of electric vertical take-off and landing (eVTOL) aircraft by pioneering companies such as Joby Aviation, EHang, Archer Aviation, and Volocopter is a critical enabler of this vision, promising a safer, quieter, and more sustainable alternative to ground-based travel [long_demand_2023]. However, the immense capital investment required to build a new network from the ground up presents a significant barrier [hae_choi_exploring_2022, ryerson_build_2014]. This challenge highlights the strategic importance of existing aviation infrastructure. For example, the New York and Los Angeles metropolitan areas contain 25 and 24 regional airports, respectively, many with surplus capacity [bonnefoy_emergence_2006, ryerson_build_2014]. Therefore, re-purposing existing and underutilized regional airports offers a viable entry point for initiating UAM operations.
However, transitioning UAM from a promising concept to a viable reality presents immense system-level challenges beyond infrastructures. Successful launch and operation of an UAM service requires solving a series of additional problems, including demand modeling and estimation, fleet sizing, network and corridor design, and real-time fleet management under uncertainty[long_demand_2023]. Among these, heterogeneous fleet composition emerges as a critical factor, as it directly influences capital investment and operating efficiency [kim_receding_2020, cao2024fleetsizespilluam]. Optimizing aircraft selection to match spatiotemporal demand variations minimizes seat wastage while ensuring service coverage, thereby crucially influencing system viability.
Furthermore, conventional transportation analysis tools are often inadequate for this task [koliogeorgi_auto-tuning_2024]. They struggle with the scale of regional analysis and often lack the ability to model the dynamic feedback loop between emerging air traffic and existing ground transportation systems. The significant computational time required by conventional simulators creates a performance bottleneck, hindering the large-scale, iterative analysis needed for robust system design and optimization [jiang_large_2024, saprykin_gemsim_2019].
@@ 2.1 UAM Demand Calculation
Strategic network planning is the cornerstone of successful implementation of UAM, and it fundamentally relies on accurate travel demand forecasting [cohen_urban_2021, long_demand_2023, rimjha_urban_2021]. This section reviews existing literature on UAM demand estimation to highlight a critical methodological limitation: the predominant use of static models that neglect the dynamic feedback between air and ground transportation systems.
Current research primarily estimates UAM demand by modeling mode choice based on the attributes of trips across different transportation options. In this framework, a trip qualifies for UAM if it provides substantial savings in travel time or cost relative to ground transportation [kim_potential_2023, hae_choi_exploring_2022, peng_hierarchical_2022] Some studies also suggest that airport access and egress times influence demand [coppola_urban_2025, pukhova_flying_2021, rimjha_urban_2021]. Researchers mainly draw on static data to perform demand analysis, including current airport location [rimjha_urban_2021, rimjha_urban_2021-1], Uber trip data [rimjha_urban_2021-1], census-level socioeconomic data [haan_are_2021, straubinger_proposing_2021], cell phone data[haan_are_2021], and ground-level transportation data [bulusu_traffic_2021, kim_potential_2023]. Different application of UAM also impact the forecasting of demand, a study in Milan found UAM airport shuttles were more attractive than on-demand air taxis, capturing a potential modal share of 2-5% for airport trips, compared to 1-3% for air taxis [coppola_urban_2025].
The critical limitation of this static approach is its inability to account for the dynamic feedback loop between UAM and ground transportation systems [jiang2024simulatingintegrationurbanair]. As UAM adoption increases, it diverts traffic from terrestrial routes, which can partially alleviate road congestion. By analyzing demand based on a fixed, pre-existing state of congestion, static models overlook this shifting equilibrium. As a result, they are likely to produce an overestimation of the long-term sustainable demand for UAM services. Future research must therefore evolve towards dynamic models that can capture the reciprocal effects between air and ground mobility to generate more realistic and reliable demand forecasts [kim_receding_2020].
@@ 2.2 UAM Operation Simulator
The operational management of UAM is exceptionally complex, involving dynamic passenger demand, intricate airspace interactions, and energy-constrained vehicle fleets [song_development_2021, thu_multivehicle_2022, 10637246]. Due to these interconnected complexities, analytical solutions are often intractable, making high-fidelity simulation platforms indispensable as virtual testbeds for developing and validating operational algorithms [chen_integrated_2024, lee_functional_2024].
@@
These simulators provide a dynamic environment to assess the performance and robustness of various management strategies before real-world deployment. For instance, in the domain of on-demand fleet scheduling, simulators are used to evaluate the effectiveness of optimization methods like Mixed-Integer Linear Programming (MILP) and imitation learning in minimizing passenger wait times [kleinbekman_rolling-horizon_2020, poddar_graph-based_2024, wu_optimal_2020]. Simulators also provide the necessary environment to assess the practical viability and effectiveness of heuristic approaches, such as Genetic Algorithms (GA) and Particle Swarm Optimization (PSO), for managing large-scale, heterogeneous fleets [kim_receding_2020].
@@
Furthermore, simulators are critical for validating advanced airspace management strategies. To ensure safety and efficiency in high-density airspace, Demand-Capacity Balancing (DCB) has emerged as another focal point, with solutions relying on advanced AI techniques like Multi-Agent Reinforcement Learning (MARL) for strategic traffic deconfliction [huang_strategic_2022, murthy_reinforcement_2025, xie_reinforcement_2021] and AI-driven urban airspace monitoring method [liu_intelligent_2025]. At the level of real-time dispatching, VertiSync, a centralized, conflict-free takeoff scheduling policy that explicitly integrates both trip request servicing and critical vehicle rebalancing within UAM networks [pooladsanj_throughput_2025]. Building on this, our previous research introduced a Multi-modal Regional-scale Traffic Simulation (MRTS), which provides an integrated environment to model the coupled dynamics of passengers, aircraft, and ground traffic, enabling more holistic algorithm validation [jiang2024simulatingintegrationurbanair, cao2024fleetsizespilluam].
@@
While these operational algorithms are powerful in theory, their practical development and large-scale validation are constrained by the performance of the simulation environments [koliogeorgi_auto-tuning_2024, saprykin_application_2022, saprykin_gemsim_2019]. For any operational strategy, a simulator must validate thousands of diverse, large-scale vehicle scenarios. As a result, the simulator’s speed directly determines the comprehensiveness and efficiency of this validation process [aqib_rapid_2019, saprykin_accelerating_2022]. Further more, this constraint is even more acute for AI-driven methods like reinforcement learning, where the simulator is not just a validation tool but an integral part of the training loop [murthy_reinforcement_2025].
@@
Presently, the vast majority of UAM simulation platforms are built on CPU-based architectures. This creates a computational bottleneck, rendering large-scale, real-time simulations computationally prohibitive and thus severely limiting the scope and speed of algorithmic validation [cao_integrating_2024]. Therefore, a critical research gap exists between the needs of advanced algorithms and the capabilities of current simulators. Our previous study has developed LPSim, a Large-Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework, which can help boost the computation process [jiang_large_2024].
@@ 3.1 UAM and Ground Trip Simulation
Our simulation’s primary objective is to seamlessly integrate UAM into existing transportation systems as a multimodal mode. This integration encompasses the entire journey, from first-mile to last-mile ground segments to the air trips. Consequently, our framework explicitly captures the intricate interplay between UAM operations and traditional ground transportation [jiang_large_2024].
@@
5:   Simulate car following, lane changing, and gap acceptance as detailed in [jiang_large_2024]
@@ 4.1 Airports and Ground Transportation Data
Urban travel demand data was obtained from the San Francisco Chained Activity Modeling Process-6 (SF-CHAMP) framework, an advanced model developed by the San Francisco County Transportation Authority (SFCTA) that predicts transportation patterns based on infrastructure, travel behaviors, and socio-economic data [outwater_san_2008, sall_evaluating_2010]. The dataset reflects a typical weekday, containing approximately 17.8 million trips over a full day and 6.6 million during the morning period (12:00 AM–12:00 PM), with each trip record detailing attributes such as origin, destination, and transportation mode [chan_simulating_2023]. Finally, the airport network consists of 21 general aviation airports in the Bay Area, with detailed specifications for each provided in Table 1.
To estimate the potential capacity of each airport in the Bay Area, we evaluated both its ground-side and air-side limits. The maximum ground transportation capacity was calculated using an average vehicle occupancy factor [united_states_department_of_transportation_federal_highway_policy__guidance_center_average_2018]. Concurrently, the maximum air-side capacity was determined by modeling continuous operations at full throughput, assuming a 90-second interval between takeoffs or landings, with the largest suitable aircraft for each facility. The resulting capacity estimates for the airport network are illustrated in Figure 3.
@@ 4.2 Aircraft’s Flight Simulation
For this study, individual flights are modeled using a detailed five-phase profile based on FAA standards [federal_aviation_administration_faa_2025]. The specific parameters for each phase are defined as follows:
@@ 4.2.3 Travel Time for Aero Trips
To determine the minimum operational interval for each airport runway, we apply the constraints outlined in FAA Order 7110.65 [federal_aviation_administration_faa_2025]. This interval is a function of two primary components: Runway Occupancy Time (ROT) and the Aircraft Separation Requirement (ASR). For eVTOLs, we assume an ROT of approximately 30 seconds, with a slightly longer ROT for general aviation aircraft like Cessna 208 Caravan. The ASR, which ensures safety between successive aircraft, is set at 60 seconds under Visual Flight Rules (VFR) for small aircraft following another. Therefore, the Minimum Time Interval (T) is the sum of these values, which yields 90 seconds ($30sROT+60sASR$). Consequently, the maximum theoretical capacity for a single runway is 40 operations per hour ($3600s/90s$), where an operation can be either a take-off or landing.